Circular No. 6524

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cb IN NGC 3510 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Tsukioka Cho, Toyama, of a supernova (mag 16.5) on unfiltered CCD frames taken on Dec. 15.707 UT with a 0.43-m f/6 reflector. SN 1996ca is located at α = 11^{h}03^{m}41^{s}.98, δ = +28^{o}54′13′′.7 (equinox 2000.0; measured using only three reference stars), which is 20′′.9 west and 65′′.7 north of the center of NGC 3510. There is no image of the supernova on Aoki’s patrol frames taken on Oct. 24 and Nov. 6, and nothing at this position on the Digital Sky Survey.

P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that P. Berlind obtained a spectrogram of SN 1996cb on Dec. 17.5 UT with the 1.5-m Tillinghast telescope; the spectrum shows this object to be a type-II supernova near maximum. Broad P-Cyg profiles of the Balmer lines dominate the spectrum. The minimum of the Hα absorption trough indicates a supernova expansion velocity of 21 000 km/s. Narrow emission lines from the host galaxy show a redshift of 770 km/s.

SCORPIUS X-1 A. G. Peele and N. E. White, Goddard Space Flight Center, NASA,

write: “The public RXTE ASM lightcurve of Sco X-1 covering the interval 1996 Feb. 22–Nov. 28 shows an intensity modulation with a period of ∼ 37 days. The period is detected using a Lomb-Scargle periodigram, and it has a false-alarm probablility of < 10^{−6}; it is also detected using minimum entropy and χ^{2}-folding techniques. Detection of the period is complicated by the well-known x-ray flaring from Sco X-1. The modulation is most evident in the low-energy (1.3–3.0-keV) channel, where the flaring amplitude is a minimum, and it is a sinusoidal-like modulation of the quiescent flux. Fitting a sine wave to the data in this channel gives an ephemeris of JD 2450160.9 (± 1.0) + (36.8 ± 0.4)E, with phase zero defined as the modulation minimum and a full amplitude of 13 ± 2 percent. The modulation amplitude for the quiescent flux decreases by a factor of 4–5 between the soft- and hard-energy (5.0–12.2-keV) channels. The 37-day modulation does not appear to be related to the flaring episodes. The modulation is not detected in archival Vela 5B and Ariel 5 ASM x-ray data (but upper limits do not rule out its presence). We encourage a search for the 37-day period in the optical and radio emission from Sco X-1 (= V818 Sco).”

1996 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green